OpenStudy (anonymous):

5 cards are drawn from a dexk of 52 cards. If all 5 are red, find the probability that they all are hearts.

OpenStudy (anonymous):

Its a conditional probability

OpenStudy (anonymous):

U know that?

OpenStudy (anonymous):

yup but i am confused about the condition that all are red...

OpenStudy (anonymous):

P(all 5 hearts / all 5 reds)

OpenStudy (anonymous):

U need to find that.

OpenStudy (anonymous):

but what will be their intersection..

OpenStudy (anonymous):

P (A/B) = P ( A and B)/P(B)

OpenStudy (anonymous):

use it. can u?

OpenStudy (anonymous):

but what will be P(A and B)

OpenStudy (anonymous):

i am confused there...

OpenStudy (anonymous):

First Find: P(all 5 are red)

OpenStudy (anonymous):

13C5..right

OpenStudy (anonymous):

So, cards are not placed back.

OpenStudy (anonymous):

And REMEMBER there are 26 red cards.

OpenStudy (anonymous):

no....they are not placed back.....i am confuzed that sample space will be 52C5 or 26C5

OpenStudy (anonymous):

see P(A | B) * (26C5)/(52C5) = ???

OpenStudy (anonymous):

First, P(all 5 are red) = 26C5/52C5

OpenStudy (anonymous):

U get that?

OpenStudy (anonymous):

okk....that i got it right above but P(A and B) will be (13C5)/(26C5) or (13C5)/52C5

OpenStudy (anonymous):

P( all 5 hearts and reds) = 13C5 / 52C5

OpenStudy (anonymous):

but the answer is not matching...

OpenStudy (anonymous):

Isnt: P(all 5 hearts / all 5 reds) = 13C5 / 26C5

OpenStudy (anonymous):

answer is 0.11

OpenStudy (anonymous):

i tried the same thing b4.....but nowhere near to the answer

OpenStudy (anonymous):

pls someone help...

OpenStudy (lgbasallote):

all 5 are hearts?? how is that possible?

OpenStudy (anonymous):

There are 13 hearts....... So it is possible.

OpenStudy (anonymous):

it is perfectly possible.....

OpenStudy (lgbasallote):

oh yeah..i was thinking something else

Parth (parthkohli):

He was thinking about suits. ;P

OpenStudy (shubhamsrg):

shouldnt it be 13C5 / 52C5 ?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

you already know you have red cards so the sample space has, as @sauravshakya said, \(\binom{26}{5}\) elements in it

OpenStudy (anonymous):

I see no mistake here: P(all 5 hearts / all 5 reds) = 13C5 / 26C5

OpenStudy (anonymous):

yes, i was saying "on no" to the previous post

OpenStudy (shubhamsrg):

oh yes yes,,wont be 52C5 there..sorry..we are taking red cards only.. so i agree with saurav's ans..

OpenStudy (anonymous):

I desperately want to say 1/6....

OpenStudy (anonymous):

5 cards are drawn from a dexk of 52 cards. If all 5 are red, find the probability that they all are hearts. 5 cards are drawn from a dexk of 52 cards, find the probability that they all are hearts. These are different, right?

OpenStudy (anonymous):

I think this question is (deliberately?) confusing. If it had said "find the CONDITIONAL probability that they all are hearts" that would have been clear. The way it is now, there is no reason to go beyond the fact that there are 5 red cards and I would say that 1/6 is a good answer. If the answer was written I would write the above, if a multiple choice it would depend on the choices. I still don't know why or how the answer could be 0.11 (1/9?)

OpenStudy (phi):

you get 0.11 if the problem is 3 cards are drawn (not 5) 13C3 / 26C3 =0.11 so it looks like a typo.